OpenStudy (anonymous):

Solve by completing the square and then use the square root method to solve. x2 - 4x - 32 = 0 A. {-8, 4} B. {-2, 16} C. {-4, 8} D. {-16, 2} Solve by completing the square and then use the square root method to solve. x2 - 4x - 32 = 0 A. {-8, 4} B. {-2, 16} C. {-4, 8} D. {-16, 2} @Mathematics

hero (hero):

x^2-4x-32 = 0 x^2-4x+8x -32 =0 x(x-4)+8(x-4)=0 (x-4)(x+8)=0 x = -8,4

hero (hero):

Are you understanding these at all?

hero (hero):

I take that as a "no"